OpenStudy (anonymous):

A submarine carries a pressure meter so that the crew can work out how far they are below the surface of the sea. At the surface, the meter indicates a pressure of 100 kPa. The density of the seawater is 1030 kg/m^3. Q.What is the depth below the surface when the meter reads 450 kPa?

OpenStudy (anonymous):

let the depth be x so at surface x= 0 P = mg / A P/x = mg / V = dg [d= density =m/V] x = P / dg = 350000 pa/ 1030 x 10 = 33.98 = 34.6 m

OpenStudy (anonymous):

Hope that helps!

OpenStudy (anonymous):

Im sorry, but i dont really understand what ure writting :(. can u explain in detail? for example ure using P = mg / A is that Pressure = mass x gravitational / ?. what is that A by the way? Area?

OpenStudy (ujjwal):

Difference in pressure, p=hdg where, p is difference in pressure h is difference in height (depth) d is density of liquid g is acceleration due to gravity